OpenStudy (anonymous):

What is the slope of the line –6y + 2x = 18 in simplest form? ? ----- ?

OpenStudy (anonymous):

@Zale101 @undeadknight26 @KendrickLamar2014

OpenStudy (zale101):

try to get the equation to become in slope intercept from, so you can visibly see the slope slope intercept form is y=mx+b where m is the slope.

OpenStudy (zale101):

–6y + 2x = 18 change it to y=mx+b that means that the equation should be set to equal to y –6y + 2x = 18 subtract 2x from both sides –6y = 18-2x divide -6 to both sides

OpenStudy (zale101):

Makes sense? @FateRocks8D , can you continue working on your problem now?

OpenStudy (anonymous):

a lil bit o.o thx @Zale101

OpenStudy (zale101):

This problem basically wants you to do algebra to solve for y

OpenStudy (anonymous):

my answer would be 1/3?

OpenStudy (zale101):

Correct. –6y = 18-2x \(\Large \frac{–6}{-6}y = \frac{18}{-6}-\frac{2}{-6}x\) \(\large y=-3+\frac{2}{6}x\) 2/6 is reduced to 1/3 --> \(\large y=\frac{1}{3}x-3\) Slope is 1/3 YAY! Good job :)

OpenStudy (zale101):

Try to refresh you page to get rid of star thingies

OpenStudy (anonymous):

Thank you :) !!! @Zale101

OpenStudy (zale101):

Sorry, i wont bring up the question. I was just recalling the methods in algebra before learning slope equations.

OpenStudy (zale101):

To make it look more simpler for you

OpenStudy (zale101):

No problem! :)

OpenStudy (anonymous):

well, thanks :D ! I could kiss u if I can xD